find all zeros of the function. f(x)x^3+72-5x^2-18x
Answer by ewatrrr(24785)


Hi
finding all zeros of the function.
f(x)= x^3-5x^2-18x +72
x^3-5x^2-18x +72 = 0 |If x=3 then zero (27 -45 -54 + 72 = 0)
(x -3)(x-6)(x+4) =0 |Dividing x^3-5x^2-18x +72 by (x-3)
(x-3)=0 x = 3
(x-6)=0 x = 6
(x+4)=0 x = -4